OpenStudy (anonymous):

5a+20 divided by a+4

OpenStudy (anonymous):

=5(a-4)/(a-4) **(a-4) reduces, which leaves with : 5

OpenStudy (anonymous):

so, it wants the restrictions of the variables so would it be \[a\]

OpenStudy (anonymous):

a cannot equal 5?

jimthompson5910 (jim_thompson5910):

what value makes the denominator a+4 equal to 0 ?

OpenStudy (anonymous):

-4?

jimthompson5910 (jim_thompson5910):

so \[\Large a \neq -4\] to avoid division by zero

OpenStudy (anonymous):

Oops I though it was reducing question

jimthompson5910 (jim_thompson5910):

that reads "a is not equal to -4"

OpenStudy (anonymous):

okay, well.. for the next one it's 2x-3² divided by 5x³ and it says that a cannot equal 0, but how do i solve the problem to get the answer?

jimthompson5910 (jim_thompson5910):

same question: what makes the denominator 0?

jimthompson5910 (jim_thompson5910):

to answer that, you set the denominator equal to zero and solve for x

OpenStudy (anonymous):

so... 1/5?

jimthompson5910 (jim_thompson5910):

no

jimthompson5910 (jim_thompson5910):

5x^3 = 0 x^3 = 0 x = cube root of 0 x = 0

jimthompson5910 (jim_thompson5910):

you can check it by plugging x =0 into 5x^3

OpenStudy (anonymous):

that's just 0 though right...?

jimthompson5910 (jim_thompson5910):

so that's why x cannot equal 0 to avoid division by zero in "2x-3² divided by 5x³ "

OpenStudy (anonymous):

but how do i solve the problem? it's either 1/5, 1/5x, 1/5x², or 2-3x/5x²

jimthompson5910 (jim_thompson5910):

oh you just want to simplify

jimthompson5910 (jim_thompson5910):

one sec

OpenStudy (anonymous):

alright, thanks

jimthompson5910 (jim_thompson5910):

The original problem is \[\Large \frac{2x - 3x^2}{5x^3}\] right?

OpenStudy (anonymous):

yes sir

jimthompson5910 (jim_thompson5910):

what can you factor out of the numerator?

OpenStudy (anonymous):

ummmmm........ 3x².?

jimthompson5910 (jim_thompson5910):

what do 2x and 3x^2 have in common? what factors?

OpenStudy (anonymous):

oh, 6?

jimthompson5910 (jim_thompson5910):

6 is the LCM of 2 and 3, I'm looking for the GCF

jimthompson5910 (jim_thompson5910):

2x has an x 3x^2 = 3*x*x So they both have 1 x That's the GCF

jimthompson5910 (jim_thompson5910):

when you factor x from the numerator, you get?

OpenStudy (anonymous):

1?

jimthompson5910 (jim_thompson5910):

2x - 3x^2 x*2 - x*3x what's the next step? We want to factor out x

OpenStudy (anonymous):

divide by 1x?

jimthompson5910 (jim_thompson5910):

yes, and have an x outside like this x(2 - 3x)

jimthompson5910 (jim_thompson5910):

if you were to distribute that x back in, you'll get 2x - 3x^2 again

jimthompson5910 (jim_thompson5910):

so we factor the numerator and denominator and then cancel to get \[\Large \frac{2x - 3x^2}{5x^3}\] \[\Large \frac{x(2 - 3x)}{5x^3}\] \[\Large \frac{x(2 - 3x)}{x*5x^2}\] \[\large \frac{\color{red}{\cancel{\color{black}{x}}}(2 - 3x)}{\color{red}{\cancel{\color{black}{x}}}*5x^2}\] \[\Large \frac{2 - 3x}{5x^2}\]

OpenStudy (anonymous):

thanks sooo much!

jimthompson5910 (jim_thompson5910):

you're welcome